# SuPreME: A Supervised Pre-training Framework for Multimodal ECG Representation Learning

### v1

@@ 5 Conclusion
We present a novel LLM-based method for ECG clinical entity extraction and introduce SuPreME, a scalable supervised pre-training framework for multimodal ECG representation learning that aligns ECG signals with fine-grained, standardized medical terminologies rather than free-text reports. Its Cardiac Fusion Network (CFN) and Clinical Knowledge-Enhanced Prompt Engineering (CKEPE) eliminate the need for further fine-tuning, enabling robust zero-shot classification with concise cardiac queries. Benchmarked on six downstream datasets, SuPreME achieves superior zero-shot performance against 11 eSSLs, underscoring both data efficiency and diagnostic precision. Our results highlight the value of explicit entity-level supervision over raw text alignment in ECG multimodal learning, providing a strong basis for clinically oriented ECG representation learning.
@@
While SuPreME demonstrates robust classification performance and superior results compared to existing eSSL frameworks, there are some limitations to consider. The framework’s reliance on LLMs for clinical entity extraction may face challenges in capturing highly specialized or ambiguous medical knowledge without domain-specific fine-tuning. Additionally, although SuPreME achieves impressive zero-shot classification performance, the approach assumes that the pre-trained model can generalize well across diverse clinical contexts, which may not always hold in real-world scenarios with varying data quality and noise levels. Future work will focus on enhancing the robustness of clinical entity extraction and developing more adaptive strategies for diverse ECG data, as well as exploring methods to further optimize the pre-training process.
Acknowledgments
This work is supported by the School of Informatics at the University of Edinburgh and the Data Science Institute at Imperial College London. We extend our gratitude to the Edinburgh Compute and Data Facility [^7] http://www.ecdf.ed.ac.uk/ (ECDF) and the Imperial College Research Computing Service [^8] http://doi.org/10.14469/hpc/2232 for providing essential computational resources for this research.
@@ Appendix A Appendix
A.1 Electrocardiogram (ECG)
@@
A.2 Implementation Details
@@
Transformer Block Structure. The Transformer architecture (Vaswani, 2017) is widely used for seq2seq modeling, learning global dependencies via self-attention instead of recurrent or convolutional structures. It consists of an encoder-decoder design, where both the encoder and decoder utilize stacked self-attention and feed-forward layers, as shown in Figure 13.
@@
Projection of Text Query Embeddings. To align the $D$-dimensional text query embeddings $E(q)$ with ECG embeddings for multimodal fusion, $\mathbf{E}(q)\in\mathbb{R}^{D}$ is projected into the same latent space $\mathbb{R}^{\text{proj\_out}}$ with ECG embeddings using a two-layer feedforward neural network with a ReLU activation function:
@@
$\mathbf{E}^{\prime}(q)=\text{Proj}(\mathbf{E}(q))=\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\cdot\mathbf{E}(q))$ (6)
where $\mathbf{W}_{1}\in\mathbb{R}^{\text{proj\_input}\times\text{hidden}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{\text{proj\_hidden}\times\text{proj\_out}}$ are the weights of the linear layers. The final embedding $\mathbf{E}^{\prime}(q)$ represents the query in the target latent space, ready for downstream multimodal fusion with ECG signal embeddings.
@@
Initialization of Cardiac Fusion Network. All weights in linear layers and attention modules are initialized with a normal distribution, $\mathbf{W}\sim\mathcal{N}(0,0.02)$. To support batch processing, the text embeddings $\mathbf{F}_{\text{text}}$ are expanded to match the batch size $B$. Both ECG and text embeddings undergo layer normalization to improve training stability and convergence.
A.3 Performance Evaluation Metric

### v2

@@ 5 Conclusion
We present a novel LLM-based method for ECG clinical entity extraction and construct a high-quality labeled dataset from MIMIC-IV-ECG. Building on this, we propose SuPreME, a scalable supervised pre-training framework for multimodal ECG representation learning that fuses ECG signals with fine-grained, standardized medical terminologies rather than free-text reports. Its Cardiac Fusion Network (CFN) and simplified Clinical Knowledge-Enhanced Prompt Engineering (CKEPE) eliminate the need for further fine-tuning, enabling robust zero-shot classification with concise cardiac queries. Benchmarked on six downstream datasets, SuPreME achieves superior zero-shot performance against 11 eSSLs, underscoring both data efficiency and diagnostic precision. Our results highlight the value of explicit entity-level supervision over raw text alignment in ECG multimodal learning, providing a strong basis for clinically oriented ECG representation learning.
@@
While SuPreME achieves strong zero-shot performance, several limitations remain. First, the clinical entity extraction relies on a large language model not fine-tuned for cardiology, which may miss rare or ambiguous terms and introduce noise. Second, SuPreME assumes generalization across clinical settings, but real-world data often involve device variability, demographic shifts, and class imbalance. Our experiments show lower performance on rare conditions, indicating sensitivity to distribution shift. Additionally, because the ECG encoder is trained jointly with CFN rather than via contrastive objectives, its features alone may not always outperform other baselines under linear probing. Lastly, most existing ECG baselines are single-modal and few of them support open zero-shot evaluation (e.g., MERL), underscoring the need for clinically motivated zero-shot benchmarks that better reflect practical deployment scenarios and support fairer comparison across methods.
@@ Appendix A Appendix
A.1 Dataset and Model Overview
A.1.1 Pre-training Dataset
MIMIC-IV-ECG. MIMIC-IV-ECG [^6] MIMIC-IV-ECG is available at https://physionet.org/content/mimic-iv-ecg/1.0/. is a comprehensive database containing 800,035 diagnostic ECG samples from 161,352 unique patients, with 12-lead recordings in 10 second length and sampled at 500 Hz Gow et al. (2023). These data have been matched with patient records in the MIMIC-IV clinical database, allowing for the association of waveforms with reports when a cardiologist’s report is available through provided linking information. To enhance the usability of the data, we exclude empty reports as well as reports containing fewer than 3 words, and replace ’NaN’ and ’Inf’ values in the ECG records with the average of 6 neighboring points. Ultimately, the dataset used for clinical entity extraction tasks includes 771,500 samples, each comprising 18 machine-generated ECG reports based on rules and the corresponding ECG data. After clinical NER and deduplication on the 18 ECG reports of each sample, the dataset holds 295 labels of professional medical terminologies.
A.1.2 Downstream Dataset
PTB-XL. PTB-XL [^7] PTB-XL is available at https://physionet.org/content/ptb-xl/1.0.3/. is a large open-source ECG dataset, comprising 21,799 clinical ECG records from 18,869 patients, with each lead sampled at a rate of 500 Hz and a duration of 10 seconds Wagner et al. (2020). A total of 71 different ECG reports are SCP-ECG compliant, covering diagnostic, form and rhythm reports. PTB-XL also provides a recommended train-test split and includes multi-level ECG annotations, covering Superclass (5 categories), Subclass (23 categories), Form (19 categories), and Rhythm (12 categories). Notably the 4 subsets have different sample sizes.
CPSC-2018. The CPSC-2018888CPSC-2018 is available at http://2018.icbeb.org/Challenge.html. dataset originates from the China Physiological Signal Challenge (CPSC) 2018, including 6,877 records from 9,458 patients, with durations ranging from 6 to 60 seconds Liu et al. (2018). The standard 12-lead ECG data is sampled at a rate of 500 Hz, collected from 11 hospitals and categorized into 9 different labels: 1 normal type and 8 abnormal types.
Chapman-Shaoxing-Ningbo (CSN). The CSN [^9] Chapman-Shaoxing-Ningbo is available at https://physionet.org/content/ecg-arrhythmia/1.0.0/. 12-lead ECG dataset is created with the support of Chapman University, Shaoxing People’s Hospital and Ningbo First Hospital, which includes 12-lead ECGs from 45,152 patients, with a sampling rate of 500 Hz and a duration of 10 seconds Zheng et al. (2020, 2022). It contains expert annotated features that cover variety of common heart rhythms and other cardiovascular conditions. We exclude ECG records with "unknown" annotations and get 23,026 ECG records with 38 different labels.
A.1.3 Llama3.1-70B-Instruct Model
Llama3.1-70B-Instruct [^10] Llama3.1-70B-Instruct is available at https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct. is a 70-billion parameter large language model released by Meta AI as part of the Llama 3 family. Built on a transformer decoder architecture, it is optimized for instruction following and few-shot generalization through extensive supervised fine-tuning and reinforcement learning from human feedback (RLHF). Compared to its predecessors, Llama3.1-70B-Instruct demonstrates substantial improvements in reasoning, factuality, and alignment with user intent across a wide range of NLP tasks.
In our framework, we leverage Llama3.1-70B-Instruct to extract fine-grained diagnostic entities from free-text ECG reports in the MIMIC-IV-ECG dataset. The scale and instruction-tuning of this model make it well suited for domain-specific named entity recognition (NER) in noisy clinical narratives. Our objective is to construct a high-quality, large-scale set of cardiac entities and their mapped terminologies, enabling robust supervision for ECG-text multimodal learning and promoting reproducibility in future research.
Although smaller models can provide acceptable results (see Section 4.4), we adopt Llama3.1-70B-Instruct to maximize annotation quality, particularly for downstream applications in clinical and low-resource settings that rely on precise structured supervision.
A.2 Electrocardiogram (ECG)
@@
A.3 Clinical NER Prompts, Statistics and Case Study
A.3.1 Prompt for Medical Database Terminology Filtering
system_message = """\
You are a clinical NLP assistant specializing in identifying ECG related terminologies from medical databases.
Your primary task is to serve as a strict terminology filter that judges whether the provided terminology is related to ECG or not, and output your judgement in a **strictly formatted JSON object** that conforms exactly to the following schema:
"IS_ECG_TERM": true/false
**Strict constraints**:
- Return **only** the JSON object. Do not include any natural language explanation or commentary.
- Do not hallucinate or invent fields not specified above.
Your output will be used in real-life clinical settings. Any deviation from this format may cause serious issues in downstream applications. Be precise and compliant.
def get_prompt(row):
return f"""\
Please read and give your judgement on the following terminology.
Terminology:
\"{row["ENG_TERM"]}\""
A.3.2 Prompt for Report Entity Extraction
system_message = """\
You are a clinical NLP assistant specializing in information extraction from medical ECG (electrocardiogram) reports. Your role is to serve as a strict, schema-aware entity extractor that produces structured annotations for downstream machine learning and clinical data analysis tasks.
Please learn the knowledge including common ECG terminologies and abbreviations first:
**Common ECG terminologies**:
Normal: "normal sinus rhythm", "normal ecg", "sinus rhythm", "within normal limits", "no abnormalities detected", ...
Abnormal: "atrial fibrillation", "ventricular tachycardia", "left ventricular hypertrophy", "right bundle branch block", "ST elevation, "T wave inversion", "prolonged QT interval", "first degree AV block", "pacemaker rhythm", ...
Uncertain: "possible infarction", "borderline ecg", "nonspecific ST-T changes", "probable left ventricular hypertrophy", "cannot rule out ischemia", ...
**Demo Abbreviations**:
NSR: "Normal Sinus Rhythm",
AFIB: "Atrial Fibrillation",
AFL: "Atrial Flutter",
VT": "Ventricular Tachycardia",
PVC: "Premature Ventricular Contraction",
PAC: "Premature Atrial Contraction",
LVH: "Left Ventricular Hypertrophy",
RVH: "Right Ventricular Hypertrophy",
RBBB: "Right Bundle Branch Block",
LBBB: "Left Bundle Branch Block",
AVB1: "First Degree AV Block",
AVB2: "Second Degree AV Block",
AVB3: "Third Degree AV Block",
STEMI: "ST-Elevation Myocardial Infarction",
NSTEMI: "Non-ST-Elevation Myocardial Infarction",
TW": "T Wave Inversion",
QTc: "Corrected QT Interval",
BBB: "Bundle Branch Block",
LAD: "Left Axis Deviation",
RAD: "Right Axis Deviation",
SA: "Sinoatrial",
PVCs: "Premature Ventricular Contractions",
PACs: "Premature Atrial Contractions"
Your primary task is to identify all relevant entities in an ECG report and then classify based on diagnosis certainty, afterwards output them in a **strictly formatted JSON object** that conforms exactly to the following schema:
‘‘‘json
"global": [...], # All ECG entities from the provided report
"classification": {
"normal": [...], # Entities confidently labeled as clinically "normal" (e.g., "normal ECG", "sinus rhythm")
"abnormal": [...], # Entities labeled as clinically "abnormal" (e.g., "atrial fibrillation", "ST elevation")
"uncertain": [...] # Entities with uncertainty or ambiguity in the report context (e.g., "possible LVH", "undetermined".)
**Strict constraints**:
- Return **only** the JSON object. Do not include any natural language explanation or commentary.
- Do not hallucinate or invent fields not specified above.
- Do not extract adjectives or modifiers (e.g., "nonspecific", "mild", "marked", "possibly", "likely") as standalone entities. If a descriptive modifier qualifies an entity (e.g., "nonspecific ST-T changes", "likely normal ecg"), include it in the full entity string.
- Do not extract entire sentences or diagnostic phrases as a single entity. If a sentence contains multiple medical concepts, extract each as a separate entity.
- If an entity contains conjunctions (e.g., "and", "or", "and/or"), causal phrases (e.g., "due to", "with"), or multiple anatomical locations (e.g., "inferior/lateral"), you must split it into separate entities.
- If there are entities with clinically same meanings in the given report, only retain one with better expression.
**Some examples**:
- [Modifier + Entity]:
Input: "lateral st-t changes are probably due to ventricular hypertrophy"
Output: {"global": ["lateral st-t changes", "ventricular hypertrophy"], "classification": {"normal": [], "abnormal": ["lateral st-t changes", "ventricular hypertrophy"], "uncertain": []}}
- [Entity A with/and/or/’/’ Entity B]:
Input: "sinus rhythm with pacs. hypertrophy and/or ischemia. inferior/lateral st-t changes."
Output: {"global": ["sinus rhythm", "pacs", "hypertrophy", "ischemia", "inferior st-t changes", "lateral st-t changes"], "classification": {"normal": ["sinus rhythm"], "abnormal": ["pacs", "hypertrophy", "ischemia", "inferior st-t changes", "lateral st-t changes"], "uncertain": []}}
- [Entity + Further Description]:
Input: "inferior infarct - age undetermined. pacemaker rhythm - no further analysis. poor r wave progression - probable normal variant."
Output: {"global": ["inferior infarct", "age undetermined", "pacemaker rhythm", "poor r wave progression", "probable normal variant"], "classification": {"normal": [], "abnormal": ["inferior infarct", "pacemaker rhythm", "poor r wave progression"], "uncertain": ["age undetermined", "probable normal variant"]}} # "no further analysis" is not a medical entity
Your output will be used in real-life clinical settings. Any deviation from this format may cause serious issues in downstream applications. Be precise and compliant.
def get_prompt(row):
return f"""\
Please extract all relevant clinical entities from the following ECG report.
Return the output strictly in the JSON format described in the system prompt.
Do not include any explanation or additional text.
ECG report text:
\"{row["total_report"]}\"
A.3.3 Case Study of Deduplication and Mapping
To address concerns about how descriptive cardiac queries are constructed and how they reduce noise compared to raw NER outputs, we present a representative case study from the MIMIC-IV-ECG dataset.
Original Clinical Report:
“Sinus rhythm w/ PACs, QTc prolonged, Left axis deviation, RBBB with left anterior fascicular block, Inferior/lateral T changes may be due to myocardial ischemia, Low QRS voltages in precordial leads.”
Extracted Raw Entities:
"sinus rhythm", "PACs", "QTc prolonged", "Left axis deviation", "RBBB", "Left anterior fascicular block", "Inferior/lateral T changes", "Myocardial ischemia", "Low QRS voltages in precordial leads"
Mapped and Standardized Queries (after Deduplication and Mapping):
SCP Code Standardized Query Matched Raw Entities (Cosine Similarity)
SR sinus rhythm sinus rhythm (1.0000)
PAC premature atrial complex PACs (0.8976)
LNGQT prolonged QT interval QTc prolonged (0.9434)
ALS axis left shift Left axis deviation (0.8723)
RBBB right bundle branch block RBBB (1.0000)
LAFB left anterior fascicular block Left anterior fascicular block (1.0000)
NT non-specific T wave changes Inferior/lateral T changes (0.7751)
MI myocardial infarction Myocardial ischemia (0.9231)
LVOLT low QRS voltages Low QRS voltages in precordial leads (0.8919)
Table 11: Example mapping from raw NER entities to standardized cardiac query labels.
This example illustrates how the same clinical concept may be expressed in different lexical forms (e.g., “PACs” vs. “premature atrial complex”) or contain verbose phrasing (e.g., “Low QRS voltages in precordial leads”), leading to noisy or redundant supervision if used directly. By clustering and mapping using MedCPT embeddings and similarity thresholds (Table 11), these expressions are unified under concise, standardized queries aligned with SCP codes.
SCP Code Standard Cardiac Query Mapped Raw NER Entities (Cosine Similarity)
NORM normal Normal (1.000), of normal (0.995), Normal result (0.979), Normal interest (0.953), percent of normal (0.953), percent of normal (0.942)
IMI inferior myocardial infarction Inferior myocardial ischemia (0.956), Inferior MI on ECG (0.935), ECG shows inferior MI (0.928), Myocardial infarction (0.923), Old inferior MI (0.907)
LVH left ventricle hypertrophy Left ventricular hypertrophy (0.992), Severe LVH (0.967), Hypertensive LVH (0.948), Acquired LVH (0.940), Congenital LVH (0.904)
Table 12: Example clusters of raw NER entities mapped to standardized cardiac queries.
In Table 12 we show parts of the clustering and deduplication results on the pre-train dataset MIMIC-IV-ECG. This process prevents redundant terms from introducing duplicate supervision, normalizes entities with modifiers (e.g., “in precordial leads”), and enforces semantic consistency across ECG samples. These standardized queries form the global label set used for training, enabling clean multimodal supervision and robust generalization in zero-shot settings.
A.4 Pre-training Framework Implementation
A.4.1 Transformer Block Structure.
@@ A.4 Pre-training Framework Implementation
The Transformer architecture (Vaswani, 2017) is widely used for seq2seq modeling, learning global dependencies via self-attention instead of recurrent or convolutional structures. It consists of an encoder-decoder design, where both the encoder and decoder utilize stacked self-attention and feed-forward layers, as shown in Figure 9.
@@
A.4.2 Projection of ECG Embeddings
Following the Transformer encoder stack in the Vision Transformer (ViT) backbone, the resulting ECG token sequence $\mathbf{Z}\in\mathbb{R}^{B\times(L\cdot N)\times D}$ is passed through a modality-specific projection head to align its dimensionality with the shared multimodal latent space used in fusion.
The projection head is implemented as a two-layer multilayer perceptron ($\text{MLP}_{\text{ECG}}$), consisting of:
  • A linear transformation from the ViT output width $D$ to an intermediate hidden size $D_{h}$;
  • A non-linear activation function (ReLU);
  • A linear transformation from $D_{h}$ to the final projected dimension $D^{\prime}$, shared with the text modality.
Formally, the projection can be written as:
@@ A.4.2 Projection of ECG Embeddings
$\begin{array}[]{c}\mathbf{Emb}_{\text{ECG}}^{\text{hidden}}=\mathbf{Z}_{\text{dropout}}\mathbf{W}_{1}+\mathbf{b}_{1}\\[5.0pt] \mathbf{Emb^{\prime}}_{\text{ECG}}^{\text{hidden}}=\text{ReLU}(\mathbf{Emb}_{\text{ECG}}^{\text{hidden}})\\[5.0pt] \mathbf{F}_{\text{ECG}}=\mathbf{Emb^{\prime}}_{\text{ECG}}^{\text{hidden}}\mathbf{W}_{2}+\mathbf{b}_{2}\end{array}$ (7)
where ReLU is the activation function, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ bias terms, and $\mathbf{W}_{1}\in\mathbb{R}^{D\times D_{h}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{D_{h}\times D^{\prime}}$ are learnable parameters.
This projection layer serves to improve non-linear representational capacity before multimodal alignment, and to map ViT-specific features to a dimensionally consistent space with text query embeddings, enabling efficient cross-modal attention in the Cardiac Fusion Network (CFN).
A.4.3 Projection of Text Query Embeddings
To align cardiac query embeddings with ECG features in the multimodal latent space, we apply a modality-specific projection head to the output of the MedCPT query encoder (QEnc). Given $M$ queries encoded into a matrix $\mathbf{E}\in\mathbb{R}^{M\times 768}$, the projection head transforms each 768-dimensional embedding into a $D^{\prime}$-dimensional representation compatible with ECG tokens.
The projection is implemented as a two-layer multilayer perceptron ($\text{MLP}_{\text{Query}}$) as well, consisting of:
  • A linear transformation from 768 to a hidden dimension $D_{h}$;
  • A non-linear activation function (GELU);
  • A linear transformation from $D_{h}$ to the target fusion dimension $D^{\prime}$.
Formally, the operation is defined as:
$\begin{array}[]{c}\mathbf{Emb}_{\text{Query}}^{\text{hidden}}=\mathbf{E}_{\text{dropout}}\mathbf{W}_{3}+\mathbf{b}_{3}\\[5.0pt] \mathbf{Emb^{\prime}}_{\text{Query}}^{\text{hidden}}=\text{GeLU}(\mathbf{Emb}_{\text{Query}}^{\text{hidden}})\\[5.0pt] \mathbf{F}_{\text{Query}}=\mathbf{Emb^{\prime}}_{\text{Query}}^{\text{hidden}}\mathbf{W}_{4}+\mathbf{b}_{4}\end{array}$ (8)
where GeLU is the activation function, $\mathbf{b}_{3}$ and $\mathbf{b}_{4}$ bias terms, and $\mathbf{W}_{3}\in\mathbb{R}^{768\times D_{h}}$ and $\mathbf{W}_{4}\in\mathbb{R}^{D_{h}\times D^{\prime}}$ are learnable parameters.
This projection head enables cross-modal alignment by transforming domain-specific textual semantics into a shared feature space used by the Cardiac Fusion Network (CFN). The structure mirrors the ECG-side projection to maintain architectural symmetry and training stability.
A.4.4 Initialization of Cardiac Fusion Network.
@@ A.4.3 Projection of Text Query Embeddings
All weights in linear layers and attention modules are initialized with a normal distribution, $\mathbf{W}\sim\mathcal{N}(0,0.02)$. To support batch processing, the text embeddings $\mathbf{F}_{\text{text}}$ are expanded to match the batch size $B$. Both ECG and text embeddings undergo layer normalization to improve training stability and convergence.
A.5 Zero-shot Evaluation Analysis
A.5.1 Specific Classification Mechanism
During zero-shot evaluation, the class set (i.e., diagnostic query set $\mathcal{Q}$) is dynamically specified per downstream dataset but remains fixed for all samples within that dataset. The model computes one score per query in $\mathcal{Q}$ for a given ECG sample. These scores are produced via a sigmoid-activated MLP head following the Cardiac Fusion Network (CFN) output, where each query representation attends over the ECG feature sequence. Importantly, this design supports variable-sized query sets across datasets, and prediction is always performed over the currently defined $\mathcal{Q}$. The classifier weights are not pre-defined or fixed, but learned representations aligned to query embeddings through cross-modal attention, ensuring full flexibility across unseen classes.
A.5.2 Simplified Clinical Knowledge-Enhanced Prompt Engineering
In our implementation of simplified CKEPE query construction, we follow the general design principle introduced in MERL Liu et al. (2024). The original CKEPE pipeline in MERL employs GPT-4 with web browsing to retrieve attributes and subtypes of each cardiac condition from clinical knowledge sources such as SNOMED CT and SCP-ECG. The prompt typically used is:
"Which attributes and subtypes does <cardiac condition> have?"
The responses are then validated against the external databases to avoid hallucination and finally organized into detailed clinical descriptions used as prompts for downstream evaluation (see MERL Section 3.4 and Figure 3).
In contrast, we adopt a simplified version of this process (Section 3.3) aimed at reducing verbosity while preserving clinical specificity. Specifically, we use GPT-4o with the following style of prompt:
"Provide the standard clinical definition of <SCP diagnostic code> based on the SCP-ECG protocol."
The generated responses are then automatically validated by external databases as well to reduce hallucinated content. Rather than expanding into all potential attributes or phenotypes (as done in MERL), we retain only the concise, high-precision diagnostic phrase for each class, enabling cleaner alignment with the downstream label space.
Take a simple case study as example, for the diagnostic class LBBB (Left Bundle Branch Block), MERL would produce a long-form prompt such as:
“A conduction abnormality characterized by delayed depolarization of the left ventricle, typically resulting in a widened QRS complex ($>$120 ms), often associated with underlying structural heart disease or ischemia.”
In contrast, our simplified prompt (after GPT-4o generation and medical verification) becomes:
“left bundle branch block”
This compact form reduces potential noise in query encoding while retaining diagnostic specificity. It aligns with our hypothesis that multimodal fusion benefits more from semantically discriminative labels than verbose natural language definitions.
A.5.3 Evaluation Metrics